# Generating multipartite nonlocality to benchmark quantum computers

### v1

@@ VII Conclusion
We have demonstrated a method to test $n$-partite Bell nonlocality on quantum computers. Quantum computer often have restricted two-qubit connectivity. We have thus pointed out that graph states are a natural choice of nonlocal states that can be readily prepared if the graph is a subgraph of the connectivity graph. Moreover, for certain graph states good Bell inequalities are known. These Bell inequalities allow for an exponential violation of the classical bound, but in turn also typically require an exponential number of measurements. On the one hand, the exponential violation makes them increasingly robust to noise. On the other hand, it is impossible to measure all terms in an experiment. We have solved this problem by proposing a method in the manner of randomized measurements, e.g., direct fidelity estimation [35, 36] or few-copy entanglement detection [37]. By sampling the terms of the Bell operator at random the number of measurements can be drastically reduced. The violation can, however, still be verified with high significance. We have gauged the significance of a result with Hoeffding’s inequality. It thus dependents on the violation observed. To assess the usefulness of the method on real devices, we have first used a simple depolarisation noise model to estimate realistic violations. Finally, we have simulated the method for the IBM Eagle quantum processor. As expected with increasing accuracy of the noise model the predicted violation shrinks. However, also the simulator of the IBM processor predicts the number of terms that have to be sampled to be much smaller than the total number of terms in the Bell inequalities.
@@ Appendix C Preparation scheme for the LC state
The connectivity of a quantum computer is a connected graph $G$. It thus has a spanning tree, i.e., a tree graph that covers all vertices of $G$. A tree graph in turn can be covered by a LC state by the following steps. At the start, all qubits are assumed to be prepared in the state |$\ket{+}^{\otimes n}$.

### v2

@@ VI Conclusion
We have demonstrated a method to test $n$-partite Bell nonlocality on quantum computers. Quantum computer often have restricted two-qubit connectivity. We have thus pointed out that graph states are a natural choice of non-local states that can be readily prepared if the graph is a subgraph of the connectivity graph. Moreover, for certain graph states good Bell inequalities are known. These Bell inequalities allow for an exponential violation of the classical bound, but, in turn, also typically require an exponential number of measurements. On the one hand, the exponential violation makes them increasingly robust to noise. On the other hand, it is impossible to measure all terms in an experiment. We have solved this problem by proposing a method in the manner of randomized measurements, e.g., direct fidelity estimation [35, 36] or few-copy entanglement detection [37]. By sampling the terms of the Bell operator at random the number of measurements can be drastically reduced. The violation can, however, still be verified with high significance. We have gauged the significance of a result with Hoeffding’s inequality. It thus dependents on the violation observed. To assess the usefulness of the method on real devices, we have first used a simple depolarisation noise model to estimate realistic violations. Finally, we have simulated the method for the IBM Eagle quantum processor. As expected with increasing accuracy of the noise model the predicted violation shrinks. However, also the simulator of the IBM processor predicts the number of terms that have to be sampled to be much smaller than the total number of terms in the Bell inequalities.
@@
Note added: After submission of our article similar ideas have been discussed in [56].
@@ Appendix C Preparation scheme for the LC state
The connectivity of a quantum computer is a connected graph $G$. It thus has a spanning tree, i.e., a tree graph that covers all vertices of $G$. In turn, a tree graph can be covered by a LC state by the following steps. At the start, all qubits are assumed to be prepared in the state |$\ket{+}^{\otimes n}$.